# On Arratia’s coupling and the Dirichlet law for the factors of a random integer

### v2

@@ Abstract.
$\mathbb{E}\,\sum_{i\geqslant 1}\big{|}\log P_{i}-V_{i}\log x\big{|}\ll 1,$
where the implied constant is absolute and $N_{x}=P_{1}P_{2}\cdots$ is the unique factorization of $N_{x}$ into primes or ones with the $P_{i}$’s being non-increasing. This establishes a conjecture of Arratia (2002), who proved that the left-hand side in the above estimate can be made $\ll\log\!\log x$. We also use this coupling to give a probabilistic proof of the Dirichlet law for the average distribution of the integer factorization into $k$ parts proved in 2023 by Leung and we improve on its error term.
@@ 1. Introduction
The distribution of the process $\mathbf{L}=(L_{1},L_{2},\ldots)$ is called the GEM distribution (of parameter 1). Lastly, we sort the components of $\mathbf{L}$ in non-increasing order to create $\mathbf{V}=(V_{1},V_{2},\ldots)$. The distribution of this process is the Poisson–Dirichlet distribution (of parameter 1) [^1] The GEM and Poisson–Dirichlet distributions have more general definitions involving typically a parameter $\theta$. In the rest of the paper, we will not be mentioning the parameter since we will always work with $\theta=1$.. We note that both $\sum_{i\geqslant 1}L_{i}$ and $\sum_{i\geqslant 1}V_{i}$ are equal to $1$ almost surely.
@@
is bounded by a positive monotone function that is $o(\log x)$ as $x\to\infty$. The random variables $N_{x}$ and $\mathbf{V}$ must be strongly correlated in this new probability space to achieve this. Indeed, if, for instance, $P_{1}$ and $V_{1}$ were independent, then we would have that $P_{1}\leqslant x^{1/3}$ and $V_{1}>2/3$ can happen at the same time with positive probability, which implies $\big{|}\log P_{1}-V_{1}\log x\big{|}\geqslant\frac{\log x}{3}$ with positive probability. Hence, a coupling with $N_{x}$ and $\mathbf{V}$ being independent (also called a trivial coupling) makes (1.1) $\gg\log x$. In 2002, Richard Arratia [2] constructed a coupling satisfying
@@
(a) Theorem 1 is optimal. Indeed, since $\log P_{i}$ can never take any value in $(0,\log 2)$, we have
$\mathbb{E}\sum_{i\geqslant 1}\big{|}\log P_{i}-V_{i}\log x\big{|}\geqslant\frac{\log 2}{3}\cdot\mathbb{E}\Big{[}\#\big{\{}i\geqslant 1:V_{i}\in[a(x),2a(x)]\big{\}}\Big{]}$
with $a(x)\coloneqq\frac{\log 2}{3\log x}$, no matter how we choose the coupling between $N_{x}$ and $\mathbf{V}$. However, the expectation above is exactly equal to $\log 2$ for any choice of $x\geqslant 2$.
(b) The way we construct $N_{x}$ inside the coupling will be with a deterministic function of some random variables. These random variables stay unchanged as $x$ grows. The coupling actually generates a random process $(N_{x})_{x\geqslant 1}$.
(c) Let $\sigma$ be a random permutation uniformly distributed in the permutation group $S_{n}$. It is well known that the factorization into primes of $N_{x}$ and the decomposition into disjoint cycles of $\sigma$ share similar statistics when $n\approx\log x$. In 2006, Arratia, Barbour and Tavaré [3] have proved that there exists a coupling between $\sigma$ and $\mathbf{V}$ such that
@@ 1.1. Application to the distribution of factorizations of random integers
Let $\omega(n)$ denote the number of distinct prime factors of $n$. If $n$ is square-free and $f\in\mathcal{F}_{k}(\bm{\alpha})$, then properties (b) and (c) of the definition above imply that
@@
We will use the class of functions $\mathcal{F}_{k}(\bm{\alpha})$ to define certain “random factorizations” into $k$ parts of a random integer. Specifically, let us fix $f\in\mathcal{F}_{k}(\bm{\alpha})$ and $x\geqslant 1$. We then define a random $k$-factorization corresponding to $f$ to be a random vector $\mathbf{D}_{x}=(D_{x,1},\dots,D_{x,k})$ taking values on $\mathbb{N}^{k}$ and satisfies the formula [^2] Strictly speaking, we only need property (a) of Definition 1.1 to define $\mathbf{D}_{x}$. But we will also need the other two properties when proving Theorem 2 below.
(1.5) $\mathbb{P}\Big{[}D_{x,i}=d_{i}\ \,\forall i\leqslant k\,\Big{|}\,N_{x}=n\Big{]}=f(d_{1},\ldots,d_{k})$
for all $k$-tuples $(d_{1},\dots,d_{k})\in\mathbb{N}^{k}$ with $d_{1}\cdots d_{k}=n$.
@@
Let $f(d_{1},\ldots,d_{k})=\tau_{k}(d_{1}\cdots d_{k})^{-1}$ with $\tau_{k}(n)$ being the number of $k$-factorizations of $n$. Then $f\in\mathcal{F}_{k}(\frac{1}{k},\ldots,\frac{1}{k})$. If $f$ is seen as a probability mass function as in (1.5), then we are sampling $D_{x,1}\cdots D_{x,k}$ uniformly among all $k$-factorizations of $N_{x}$.
@@
Let $f(d_{1},\ldots,d_{k})=\prod_{j=1}^{k-1}\tau(d_{j}\cdots d_{k})^{-1}$ with $\tau(n)$ being the number of divisors of $n$. Then $f\in\mathcal{F}_{k}(\frac{1}{2},\frac{1}{4},\ldots,\frac{1}{2^{k-1}},\frac{1}{2^{k-1}})$. One way to realize this random $k$-factorization is by first sampling uniformly an divisor $D_{x,1}$ of $N_{x}$. Then, for all $j<k$, we recursively sample $D_{x,j}$ uniformly among the divisors of $\frac{N_{x}}{D_{x,1}\cdots D_{x,j-1}}$.
@@
For any fixed $\bm{\alpha}\in\Delta^{k-1}$, let $f(d_{1},\ldots,d_{k})\coloneqq\prod_{i=1}^{k}\alpha_{i}^{\Omega(d_{i})}\cdot\prod_{p|n}\binom{\nu_{p}(d_{1}\cdots d_{k})}{\nu_{p}(d_{1}),\ldots,v_{p}(d_{k})}$ with $\nu_{p}(d)$ being the $p$-valuation of $d$ and $\Omega(d)$ being the number of prime factors of $d$ counted with multiplicity. This function $f$ is in $\mathcal{F}_{k}(\bm{\alpha})$. This sampling can be understood as considering a sequence of i.i.d. random variables $(B_{i})_{i\geqslant 1}$ satisfying $\mathbb{P}[B_{i}=j]=\alpha_{j}$ and constructing the $k$-factorization $D_{x,1}\cdots D_{x,k}=N_{x}$ as
$D_{x,j}\coloneqq\prod_{\begin{subarray}{c}i\geqslant 1:\ B_{i}=j\end{subarray}}P_{i},$
@@
where $P_{1}P_{2}\cdots$ is the prime factorization of $N_{x}$ as before. With this definition, the vectors $(\nu_{p}(D_{x,1}),\cdots,\nu_{p}(D_{x,k}))$ all follow multinomial distributions for every prime $p$.
When $k=2$, we write for simplicity $D_{x}=D_{x,1}$. In Examples 1 and 2, the samplings are exactly the same when $k=2$. In both cases, we are sampling uniformly a divisor $D_{x}$ of $N_{x}$. In 1979, Deshouillers, Dress and Tenenbaum [7] proved that
(1.6) $\mathbb{P}\big{[}D_{x}\leqslant N_{x}^{u}\big{]}=\frac{2}{\pi}\arcsin\sqrt{u}+O\Big{(}(\log x)^{-\frac{1}{2}}\Big{)}$
uniformly for $x\geqslant 2$ and $u\in[0,1]$. Their proof uses crucially the Landau–Selberg–Delange method to compute sums of multiplicative functions. Their result is optimal if we want an error term uniform in $u\in[0,1]$.
It turns out that for general values of $k$ and $\bm{\alpha}$, the distribution of $\mathbf{D}_{x}$ converges to the Dirichlet distribution $\text{Dir}(\bm{\alpha})$. Recall that if $\bm{\alpha}\in\Delta^{k-1}$ satisfies $\alpha_{i}>0$ for all $i$, we say that a $\Delta^{k-1}$-valued random vector $\mathbf{Z}$ follows the Dirichlet distribution $\mathrm{Dir}(\bm{\alpha})$ if
@@
$\mathbb{P}\big{[}Z_{i}\leqslant u_{i}\ \,\forall i<k]$ $=F_{\bm{\alpha}}(\mathbf{u})\coloneqq\prod_{i=1}^{k}\Gamma(\alpha_{i})^{-1}\mathop{\int\cdots\int}\limits_{\begin{subarray}{c}0\leqslant t_{i}\leqslant u_{i}\ \forall i<k\\ t_{1}+\cdots+t_{k-1}\leqslant 1\end{subarray}}\prod_{i=1}^{k}t_{i}^{\alpha_{i}-1}\,\mathrm{d}t_{1}\cdots\,\mathrm{d}t_{k-1}$
with $t_{k}\coloneqq 1-(t_{1}+\cdots+t_{k-1})$ in the integrand, for any $\mathbf{u}\in[0,1]^{k-1}$.
In 2007, Bareikis and Manstavičius [4] proved that if $f\in\mathcal{F}_{2}(\alpha_{1},\alpha_{2})$ with $\alpha_{1},\alpha_{2}>0$ and $\alpha_{1}+\alpha_{2}=1$, we have
$\mathbb{P}\big{[}D_{x}\leqslant N_{x}^{u}\big{]}=F_{(\alpha_{1},\alpha_{2})}(u)+O_{\alpha_{1},\alpha_{2}}\Big{(}(1+u\log x)^{-\alpha_{1}}(1+(1-u)\log x)^{-\alpha_{2}}\Big{)}.$
As a matter of fact, their result covered a more general class of functions than $\mathcal{F}_{2}(\alpha_{1},\alpha_{2})$, with $f(p,1)$ being allowed to be on average $\alpha_{1}$, instead of being fixed.
In 2013, Nyandwi and Smati [15] considered the $3$-factorization $D_{x,1}D_{x,2}D_{x,3}=N_{x}$ with its distribution $f$ as in Example 1. They proved
$\mathbb{P}\big{[}D_{x,1}\leqslant N_{x}^{u_{1}}\text{ and }D_{x,2}\leqslant N_{x}^{u_{2}}\big{]}=F_{(\frac{1}{3},\frac{1}{3},\frac{1}{3})}(u_{1},u_{2})+O\Big{(}(\log x)^{-\frac{1}{3}}\Big{)}.$
In 2016, de la Bretèche and Tenenbaum [6] considered the $3$-factorization $D_{x,1}D_{x,2}D_{x,3}=N_{x}$ with its distribution $f$ as in Example 2, and they found that
$\mathbb{P}\big{[}D_{x,1}\leqslant N_{x}^{u_{1}}\text{ and }D_{x,2}\leqslant N_{x}^{u_{2}}\big{]}=F_{(\frac{1}{2},\frac{1}{4},\frac{1}{4})}(u_{1},u_{2})+O\Big{(}(\log x)^{-\frac{1}{4}}\Big{)}.$
More generally, for each fixed $k\geqslant 2$ and each $\bm{\alpha}\in\Delta^{k-1}$, Leung [14, Theorem 7.1] proved in 2023 that for any $f\in\mathcal{F}_{k}(\bm{\alpha})$, we have
(1.7) $\mathbb{P}\big{[}D_{x,i}\leqslant N_{x}^{u_{i}}\ \,\forall i<k\big{]}=F_{\bm{\alpha}}(\mathbf{u})+O_{\bm{\alpha}}\Big{(}(\log x)^{-\min\{\alpha_{1},\ldots,\alpha_{k}\}}\Big{)}$
uniformly for $x\geqslant 2$ and $\mathbf{u}\in[0,1]^{k-1}$ satisfying $u_{1}+\cdots+u_{k-1}\leqslant 1$. Similarly to the result of Bareikis and Manstavičius in [4], Leung’s theorem holds for a more general class of functions than $\mathcal{F}_{k}(\bm{\alpha})$, where the quantities $f(1,\ldots,1,p,1,\ldots,1)$, with $p$ being at the $j$th coordinate, are allowed to equal $\alpha_{j}$ on average instead of pointwise.
@@
All results mentioned above use Fourier-analytic techniques, such as the Landau–Selberg–Delange method, as their main ingredients to get to their results.
@@
We have a new approach to this problem: Since the size of the divisors on a logarithmic scale of any integer is entirely determined by the size of its prime factors, we might expect Leung’s theorem to be a consequence of some quantitative form of Billingsley’s Theorem. Indeed, using the coupling from Theorem 1, we deduce an improved form of Leung’s theorem for the class of functions $\mathcal{F}_{k}(\bm{\alpha})$.
@@
Let $k\geqslant 2$, let $\bm{\alpha}\in\Delta^{k-1}$ be fixed with $\alpha_{i}>0$ for all $i$, and let $f\in\mathcal{F}_{k}(\bm{\alpha})$. In addition, let $x>1$ and let $\mathbf{D}_{x}$ be the random $k$-factorization corresponding to $f$.
@@
$\mathbb{P}\big{[}D_{x,i}\leqslant N_{x}^{u_{i}}\ \,\forall i<k\big{]}=F_{\bm{\alpha}}(\mathbf{u})+O\Bigg{(}\sum_{\begin{subarray}{c}1\leqslant i<k\\ u_{i}\neq 1\end{subarray}}\frac{1}{(1+u_{i}\log x)^{1-\alpha_{i}}(1+(1-u_{i})\log x)^{\alpha_{i}}}\Bigg{)}\,;$

### v3

@@ Abstract.
$\mathbb{E}\,\sum_{i\geqslant 1}\big{|}\log P_{i}-V_{i}\log x\big{|}\asymp 1,$
where the implied constants are absolute and $N_{x}=P_{1}P_{2}\cdots$ is the unique factorization of $N_{x}$ into primes or ones with the $P_{i}$’s being non-increasing. This establishes a 2002 conjecture of Arratia, who constructed a coupling for which the left-hand side in the above estimate is $\ll\log\!\log x$, and who also proved that the left-hand side is $\geqslant 1-o(1)$ for all couplings. In addition, we use our refined coupling to give a probabilistic proof of the Dirichlet law for the average distribution of the integer factorization into $k$ parts proved in 2023 by Leung and we improve on its error term.
@@ 1. Introduction
The distribution of the process $\mathbf{L}=(L_{1},L_{2},\ldots)$ is called the GEM distribution (of parameter 1), named after Griffiths, Engen and McCloskey (see [10, Chapter 41] for a discussion on the history of this distribution). Lastly, we sort the components of $\mathbf{L}$ in non-increasing order to create $\mathbf{V}=(V_{1},V_{2},\ldots)$. The distribution of this process is the Poisson–Dirichlet distribution (of parameter 1) [^1] The GEM and Poisson–Dirichlet distributions have more general definitions involving typically a parameter $\theta$. In the rest of the paper, we will not be mentioning the parameter since we will always work with $\theta=1$.. We note that both $\sum_{i\geqslant 1}L_{i}$ and $\sum_{i\geqslant 1}V_{i}$ are equal to $1$ almost surely.
@@
is bounded by a positive monotone function that is $o(\log x)$ as $x\to\infty$. The random variables $N_{x}$ and $\mathbf{V}$ must be strongly correlated in this new probability space to achieve this. Indeed, if, for instance, $P_{1}$ and $V_{1}$ were independent, then we would have that $P_{1}\leqslant x^{1/3}$ and $V_{1}>2/3$ can happen at the same time with positive probability, which implies $\big{|}\log P_{1}-V_{1}\log x\big{|}\geqslant\frac{\log x}{3}$ with positive probability. Hence, a coupling with $N_{x}$ and $\mathbf{V}$ being independent (also called a trivial coupling) makes (1.1) $\asymp\log x$.
In 2002, Richard Arratia [2] constructed a coupling satisfying
@@
(a) Theorem 1 is optimal in the sense that no coupling of $N_{x}$ and $\mathbf{V}$ can make this expectation tend to 00 as $x\to\infty$. This follows directly from the triangle inequality:
$\mathbb{E}\sum_{i\geqslant 1}\big{|}\log P_{i}-V_{i}\log x\big{|}\geqslant\mathbb{E}\Big{|}\sum_{i\geqslant 1}\big{(}\log P_{i}-V_{i}\log x\big{)}\Big{|}=\mathbb{E}\big{[}\log(x/N_{x})\big{]}=1-o(1).$
(b) If $X$ and $Y$ are two random variables taking values in a separable metric space $(S,\mathrm{d})$, their Wasserstein distance equals
(1.3) $\mathrm{d}_{\mathrm{W}}(X,Y)\coloneqq\inf\mathbb{E}\big{[}\mathrm{d}(X,Y)\big{]},$
where the infimum is taken over all possible couplings of $X$ and $Y$. Taking $S$ to be the $\ell^{1}$ space of real valued sequences with its associated metric, and using remark (a) above, we find that Theorem 1 is equivalent to the statement that $\mathrm{d}_{\mathrm{W}}(\mathbf{P}_{x},\mathbf{V})\asymp 1/\log x$, where $\mathbf{P}_{x}\coloneqq\big{(}\frac{\log P_{1}}{\log x},\frac{\log P_{2}}{\log x},\ldots\big{)}$.
(c) The coupling we construct contains a copy of the random process $(N_{x})_{x\geqslant 1}$. This fact might be useful in applications of Theorem 1.
(d) Let $\sigma$ be a random permutation uniformly distributed in the permutation group $S_{n}$. It is well known that the factorization into primes of $N_{x}$ and the decomposition into disjoint cycles of $\sigma$ share similar statistics when $n\approx\log x$. In 2006, Arratia, Barbour and Tavaré [3] have proved that there exists a coupling between $\sigma$ and $\mathbf{V}$ such that
@@ 1.1. Application to the distribution of factorizations of random integers
We now give some applications of Theorem 1 to the theory of divisors of integers.
Let $\Delta^{\infty}$ be the set of sequences $(v_{i})_{i\geqslant 1}$ of non-negative reals satisfying $\sum_{i\geqslant 1}v_{i}=1$, and let $\psi\colon\Delta^{\infty}\to\mathbb{R}$ be a $K$-Lipschitz function with respect to the $\ell^{1}$-metric on $\Delta^{\infty}$. Since the diameter of $\Delta^{\infty}$ is $2$, the image of $\psi$ is contained in an interval of length $2K$. Consequently, $\psi$ is bounded, and we have $\mathbb{E}[\mathbbm{1}_{N_{x}=1}\cdot\psi(\mathbf{V})]\ll_{\psi}1/x$. By Theorem 1, it follows that
(1.5) $\mathbb{E}\Big{[}\mathbbm{1}_{N_{x}\geqslant 2}\cdot\psi\big{(}\tfrac{\log P_{1}}{\log N_{x}},\tfrac{\log P_{2}}{\log N_{x}},\ldots\big{)}\Big{]}=\mathbb{E}\Big{[}\psi\big{(}V_{1},V_{2},\ldots\big{)}\Big{]}+O_{\psi}(1/\log x).$
Using this simple fact, we obtain:
Corollary 1.1.
Let $\rho(n)\coloneqq\min\{d|n:d\geqslant\sqrt{n}\}$. Then there is a constant $c\in(1/2,1)$ such that
$\sum_{n\leqslant x}\log\rho(n)=cx\log x+O(x)\qquad(x\geqslant 1).$
Proof.
Let $\psi(v_{1},v_{2},\ldots)\coloneqq\inf\big{(}\mathcal{D}(v_{1},v_{2},\ldots)\cap[\frac{1}{2},1]\big{)}$ with $\mathcal{D}(v_{1},v_{2},\ldots)\coloneqq\{\sum_{i\in B}v_{i}:B\subseteq\mathbb{N}\}$. It is easy to check that this is a $1$-Lipschitz function. Note that the left-hand side of (1.5) with this definition for $\psi$ is exactly equal to $\frac{1}{\left\lfloor x\right\rfloor}\sum_{2\leqslant n\leqslant x}\frac{\log\rho(n)}{\log n}$. With (1.5) and partial summation, we prove the corollary with $c=\mathbb{E}[\psi(\mathbf{V})]$. ∎
For the next application, we use the coupling again to extract information about integer factorizations. However, to obtain the desired result, we require much more than just the expectation in Theorem 1.
@@
Let $\omega(n)$ denote the number of distinct prime factors of $n$. If $d_{1}\cdots d_{k}$ is a square-free number and $f\in\mathcal{F}_{k}(\bm{\alpha})$, then properties (b) and (c) of the definition above imply that
@@
We will use the class of functions $\mathcal{F}_{k}(\bm{\alpha})$ to define certain “random factorizations” into $k$ parts of a random integer. Specifically, let us fix $f\in\mathcal{F}_{k}(\bm{\alpha})$ and $x\geqslant 1$. We then define a random $k$-factorization corresponding to $f$ to be a random vector $\mathbf{D}_{f,x}=(D_{f,x,1},\dots,D_{f,x,k})$ taking values on $\mathbb{N}^{k}$ and satisfying the formula [^2] Strictly speaking, we only need property (a) of Definition 1.2 to define $\mathbf{D}_{f,x}$. But we will also need the other two properties when proving Theorem 2 below.
(1.7) $\mathbb{P}\Big{[}D_{f,x,i}=d_{i}\ \,\forall i\leqslant k\,\Big{|}\,N_{x}=n\Big{]}=f(d_{1},\ldots,d_{k})$
for all $n\in\mathbb{N}$ and all $k$-tuples $(d_{1},\dots,d_{k})\in\mathbb{N}^{k}$ with $d_{1}\cdots d_{k}=n$.
@@
Let $f_{\mathrm{U}(k)}(d_{1},\ldots,d_{k})\coloneqq\tau_{k}(d_{1}\cdots d_{k})^{-1}$ with $\tau_{k}(n)$ being the number of $k$-factorizations of $n$. Then $f_{\mathrm{U}(k)}\in\mathcal{F}_{k}(\frac{1}{k},\ldots,\frac{1}{k})$. If $f_{\mathrm{U}(k)}$ is seen as a probability mass function as in (1.7), then we are sampling $D_{f_{\mathrm{U}(k)},x,1}\cdots D_{f_{\mathrm{U}(k)},x,k}$ uniformly among all $k$-factorizations of $N_{x}$.
@@
Let $f_{\mathrm{R}(k)}(d_{1},\ldots,d_{k})\coloneqq\prod_{j=1}^{k-1}\tau(d_{j}\cdots d_{k})^{-1}$ with $\tau(n)$ being the number of divisors of $n$. Then $f_{\mathrm{R}(k)}\in\mathcal{F}_{k}(\frac{1}{2},\frac{1}{4},\ldots,\frac{1}{2^{k-1}},\frac{1}{2^{k-1}})$. One way to realize this random $k$-factorization is by first sampling uniformly a divisor $D_{f_{\mathrm{R}(k)},x,1}$ of $N_{x}$. Then, for all $j<k$, we recursively sample $D_{f_{\mathrm{R}(k)},x,j}$ uniformly among the divisors of $\frac{N_{x}}{D_{f_{\mathrm{R}(k)},x,1}\cdots D_{f_{\mathrm{R}(k)},x,j-1}}$.
@@
For any fixed $\bm{\alpha}\in\Delta^{k-1}$, let $f_{\mathrm{M}(\bm{\alpha})}(d_{1},\ldots,d_{k})\coloneqq\prod_{i=1}^{k}\alpha_{i}^{\Omega(d_{i})}\cdot\prod_{p|n}\binom{\nu_{p}(d_{1}\cdots d_{k})}{\nu_{p}(d_{1}),\ldots,\nu_{p}(d_{k})}$ with $\nu_{p}(d)$ being the $p$-valuation of $d$ and $\Omega(d)$ being the number of prime factors of $d$ counted with multiplicity. The function $f_{\mathrm{M}(\bm{\alpha})}$ is in $\mathcal{F}_{k}(\bm{\alpha})$. This sampling can be understood as considering a sequence of i.i.d. random variables $(B_{i})_{i\geqslant 1}$ satisfying $\mathbb{P}[B_{i}=j]=\alpha_{j}$ and constructing the $k$-factorization $D_{f_{\mathrm{M}(\bm{\alpha})},x,1}\cdots D_{f_{\mathrm{M}(\bm{\alpha})},x,k}=N_{x}$ as
$D_{f_{\mathrm{M}(\bm{\alpha})},x,j}\coloneqq\prod_{\begin{subarray}{c}i\geqslant 1:\ B_{i}=j\end{subarray}}P_{i},$
@@
where $P_{1}P_{2}\cdots$ is the prime factorization of $N_{x}$ as before. With this definition, the vectors $(\nu_{p}(D_{f_{\mathrm{M}(\bm{\alpha})},x,1}),\cdots,\nu_{p}(D_{f_{\mathrm{M}(\bm{\alpha})},x,k}))$ all follow multinomial distributions for every prime $p$.
Starting with the pioneering work of Deshouillers, Dress, and Tenenbaum [7], many different people have studied the distribution of $\mathbf{D}_{f,x}$ for various choices of $f\in\mathcal{F}_{k}(\bm{\alpha})$, and they proved that it converges on a logarithmic scale to a certain Dirichlet distribution. Recall that if $\bm{\alpha}\in\Delta^{k-1}$ satisfies $\alpha_{i}>0$ for all $i$, we say that a $\Delta^{k-1}$-valued random vector $\mathbf{Z}$ follows the Dirichlet distribution $\operatorname{Dir}(\bm{\alpha})$ if
@@
$\mathbb{P}\big{[}Z_{i}\leqslant u_{i}\ \,\forall i<k\big{]}=F_{\bm{\alpha}}(\mathbf{u})\coloneqq\prod_{i=1}^{k}\Gamma(\alpha_{i})^{-1}\mathop{\int\cdots\int}\limits_{\begin{subarray}{c}0\leqslant t_{i}\leqslant u_{i}\ \forall i<k\\ t_{1}+\cdots+t_{k-1}\leqslant 1\end{subarray}}\prod_{i=1}^{k}t_{i}^{\alpha_{i}-1}\,\mathrm{d}t_{1}\cdots\,\mathrm{d}t_{k-1}$
With this notation, most known results on $\mathbf{D}_{f,x}$ can be stated in the following form: uniformly for $x\geqslant 2$ and $\mathbf{u}\in[0,1]^{k-1}$, we have
(1.8) $\mathbb{P}\big{[}D_{f,x,i}\leqslant N_{x}^{u_{i}}\ \,\forall i<k\big{]}=F_{\bm{\alpha}}(\mathbf{u})+O_{\bm{\alpha}}\Big{(}(\log x)^{-\min\{\alpha_{1},\ldots,\alpha_{k}\}}\Big{)}$
As we indicated above, the first result of this type was obtained in 1979 by Deshouillers, Dress, and Tenenbaum [7], who proved (1.8) for $f=f_{\mathrm{U}(2)}$ (which is also equal to $f_{\mathrm{R}(2)}$; both cases involve sampling a divisor of $N_{x}$ uniformly). For higher values of $k$, Nyandwi and Smati [19] extended this result in 2013 by proving (1.8) for $f=f_{\mathrm{U}(3)}$, and de la Bretèche and Tenenbaum [6] further proved it for $f=f_{\mathrm{R}(3)}$ in 2016. In 2023, Leung [16] demonstrated that (1.8) applies to $f=f_{\mathrm{U}(k)}$ for all $k\geqslant 2$. Leung also claimed that his proof could be adapted to any $f$ in a more general class of $k$-dimensional multiplicative functions than $\mathcal{F}_{k}(\bm{\alpha})$ (for a proof, see [17]). This more general class allows the quantities $f(1,\ldots,1,p,1,\ldots,1)$, with $p$ being at the $j$th coordinate, to equal $\alpha_{j}$ on average rather than pointwise.
Returning to 2007, Bareikis and Manstavičius [4] generalized the result of Deshouillers, Dress, and Tenenbaum with an improvement on the error term. They showed that if $f$ belongs to the more general class of two-dimensional multiplicative functions containing $\mathcal{F}_{2}(\alpha_{1},\alpha_{2})$, as described in Leung’s claim, then (1.8) holds with the improved error term
$\mathbb{P}\big{[}D_{f,x,1}\leqslant N_{x}^{u}\big{]}=F_{(\alpha_{1},\alpha_{2})}(u)+O_{\alpha_{1},\alpha_{2}}\Big{(}(1+u\log x)^{-\alpha_{1}}(1+(1-u)\log x)^{-\alpha_{2}}\Big{)}.$
@@
All results mentioned above rely on Fourier-analytic techniques, such as the Landau–Selberg–Delange method, as their main ingredients to get to their results.
@@
We have a new approach to this problem: since the size of the divisors on a logarithmic scale of any integer is entirely determined by the size of its prime factors, one might expect that the distribution of $\mathbf{D}_{f,x}$ can be accessed via a quantitative form of Billingsley’s theorem. Indeed, using the coupling from Theorem 1, we establish the following result, which proves (1.8) in full generality with a Bareikis-Manstavičius type error-term.
@@
Let $k\geqslant 2$, let $\bm{\alpha}\in\Delta^{k-1}$ be fixed with $\alpha_{i}>0$ for all $i$, and let $f\in\mathcal{F}_{k}(\bm{\alpha})$. In addition, let $x>1$ and let $\mathbf{D}_{f,x}$ be the random $k$-factorization corresponding to $f$.
@@
$\mathbb{P}\big{[}D_{f,x,i}\leqslant N_{x}^{u_{i}}\ \,\forall i<k\big{]}=F_{\bm{\alpha}}(\mathbf{u})+O\Bigg{(}\sum_{\begin{subarray}{c}1\leqslant i<k\\ u_{i}\neq 1\end{subarray}}\frac{1}{(1+u_{i}\log x)^{1-\alpha_{i}}(1+(1-u_{i})\log x)^{\alpha_{i}}}\Bigg{)}\,;$
@@
Remarks.
(a) When $u\in[0,1/2]$, we have $(1+u_{i}\log x)^{1-\alpha_{i}}(1+(1-u_{i})\log x)^{\alpha_{i}}\geqslant(0.5\log x)^{\alpha_{i}}$. Similarly, when $u\in[1/2,1]$, we have $(1+u_{i}\log x)^{1-\alpha_{i}}(1+(1-u_{i})\log x)^{\alpha_{i}}\geqslant(0.5\log x)^{1-\alpha_{i}}$. We thus find that the expression in the big-Oh in Theorem 2 is $\leqslant(0.5\log x)^{-\min\{\alpha_{1},\dots,\alpha_{k},1-\alpha_{1},\dots,1-\alpha_{k}\}}$ uniformly in $\mathbf{u}\in[0,1]^{k-1}$. Since $1-\alpha_{j}\geqslant\alpha_{i}$ for all $i\neq j$ by our assumption that $\alpha_{1}+\cdots+\alpha_{k}=1$, we conclude that the error term in Theorem 2 is $\ll_{\bm{\alpha}}(\log x)^{-\min\{\alpha_{1},\dots,\alpha_{k}\}}$, thus recovering Leung’s estimate (1.8) when $f$ lies in the class $\mathcal{F}_{k}(\bm{\alpha})$.
(b) If we relied solely on Arratia’s original coupling from [2], which leads to (1.2), our approach would yield the same Theorem 2 with each instance of $\log x$ in the error term being replaced by $\frac{\log x}{\log_{2}x}$. This adjustment would still offer an improvement over Leung’s result if all $u_{i}$ values are sufficiently far from 00 or $1$. We provide further explanations about this point in the remarks following Lemma 9.4.
The proof of Theorem 2 is based on a 1987 result of Donnelly and Tavaré [8], who proved the following probabilistic version of Leung’s theorem: If $\mathbf{V}=(V_{1},V_{2},\ldots)$ is a Poisson–Dirichlet process, and $(C_{i})_{i\geqslant 1}$ is a sequence of i.i.d. random variables that is independent of $\mathbf{V}$, supported on $\{1,\ldots,k\}$ and satisfying $\mathbb{P}[C_{i}=j]=\alpha_{j}$ for all $j$, then
(1.9) $\Bigg{(}\sum_{i:\ C_{i}=1}V_{i},\ldots,\sum_{i:\ C_{i}=k}V_{i}\Bigg{)}$
follows exactly the Dirichlet distribution $\operatorname{Dir}(\bm{\alpha})$. In 1998, Arratia [1] used this result to show that $\mathbb{P}\big{[}D_{f_{\mathrm{U}(2)},x,1}\leqslant N_{x}^{u}\big{]}$ is $F_{(\frac{1}{2},\frac{1}{2})}(u)+o(1)$ as $x\to\infty$ with probabilistic methods. We use the coupling to bridge between the distribution of (1.9) and Theorem 2 and get an explicit error term.